OpenStudy (sky7133):

Help please. At a bake sale brownies were sold for $0.75 each and cookies were sold for $0.50 each. There are 10 more cookies sold than brownies, and the total amount earned was $40. How many brownies and cookies were sold at the bake sale?

OpenStudy (trojanpoem):

P(B) = 0.75 $ P(C) = 0.50 $ let N(B) be number of brownies meanwhile, N(C) points to number of cookies N(C)_sold = N(B)_sold + 10 N(C)_sold * P(C) + N(B)_sold * P(B)= 40 (N(B)_sold + 10) * 0.50 + N(B)_sold * 0.75 = 40 0.5N(B)_sold + 0.75N(B)_sold = 40 - 5 1.25 N(B)_sold = 35 N(B)_sold = 28 brownie N(C)_sold = 38 cookie

OpenStudy (sky7133):

Thanks!